# Hypercube-RAG: Hypercube-Based Retrieval-Augmented Generation for In-domain Scientific Question-Answering

### v1

@@ Abstract
Large language models (LLMs) often need to incorporate external knowledge to solve theme-specific problems. Retrieval-augmented generation (RAG), which empowers LLMs to generate more qualified responses with retrieved external data and knowledge, has shown its high promise. However, traditional semantic similarity-based RAGs struggle to return concise yet highly relevant information for domain knowledge-intensive tasks, such as scientific question-answering (QA). Built on a multi-dimensional (cube) structure called Hypercube, which can index documents in an application-driven, human-defined, multi-dimensional space, we introduce the Hypercube-RAG, a novel RAG framework for precise and efficient retrieval. Given a query, Hypercube-RAG first decomposes it based on its entities and topics and then retrieves relevant documents from cubes by aligning these decomposed components with hypercube dimensions. Experiments on three in-domain scientific QA datasets demonstrate that our method improves accuracy by and boosts retrieval efficiency by , measured as relative gains over the strongest RAG baseline. More importantly, our Hypercube-RAG inherently offers explainability by revealing the underlying predefined hypercube dimensions used for retrieval. The code and data sets are available at https://github.com/JimengShi/Hypercube-RAG.
Hypercube-RAG: Hypercube-Based Retrieval-Augmented Generation
for In-domain Scientific Question-Answering
Jimeng Shi1, Sizhe Zhou2, Bowen Jin2, Wei Hu2, Shaowen Wang2, Giri Narasimhan1, Jiawei Han2 1Florida International University 2University of Illinois Urbana-Champaign {jshi008,giri}@fiu.edu, {sizhez,bowenj4,weih9,shaowen,hanj}@illinois.edu
1 Introduction
@@
Large language models (LLMs) often suffer from hallucinations and factual inaccuracies, especially for in-domain scientific question-answering. To address this phenomenon, retrieval-augmented generation (RAG) has emerged as the de facto approach Huang et al. (2025), incorporating external domain knowledge to generate contextually relevant responses Huang et al. (2025). Despite recent advancements, they exhibit key limitations in theme-specific applications on which the fine-grained topics are highly focused. For instance, retrieved documents are often semantically similar to the query but overlook specialized themes (e.g., terminology or nuanced contextual cues) prevalent in specific literature or reports Ding et al. (2024). Moreover, the retrieval process often suffers from inefficiency and limited transparency, which are critical in high-stakes domains, such as environment Wang et al. (2025), traffic Hussien et al. (2025), and healthcare Pujari et al. (2023).
@@
The conventional sparse lexical retrieval ranks a set of documents based on the appearance of query terms Kadhim (2019). Due to its reliance on the exact token overlap, it precisely captures specific themes or topics, and the retrieval process is efficient and interpretable. However, it struggles to retrieve contextually related or paraphrased documents that require semantic understanding. To mitigate the above limitation, dense embedding retrievers were proposed by computing the similarity between vector embeddings of query-document pairs Karpukhin et al. (2020). While contextual understanding is relatively improved, the hurdle of missing specific themes remains, causing semantically similar but off-topic retrievals Krishna et al. (2021); Kang et al. (2024). Figure 1 shows an example where the Semantic-RAG returns documents related to precipitation in Florida, but overlooks the critical location-specific information for ‘‘Melbourne Beach’’. Moreover, retrieval in an embedding space makes it challenging to interpret why certain documents were selected Ji et al. (2019). Additionally, a graph-based RAG represents documents as graph structures, where nodes correspond to entities or concepts and edges capture their relationships Peng et al. (2024). They offer explanations by outputting the traversed subgraph. However, traversing such subgraphs introduces a significant risk of information overload, as the expanding neighborhoods may include substantial amounts of irrelevant themes Wang et al. (2024b). Graph traversal also poses significant scalability bottlenecks due to its inefficient computational complexity of , where and are the number of nodes and edges in the traversed subgraph Devezas and Nunes (2021).
@@
The methods discussed above tend to perform poorly in at least one of the aspects among accuracy, efficiency, and explainability. This motivates us to develop a RAG system that overcomes this weakness. In this pursuit, we identified the text cube as a promising technique Tao et al. (2018); Wang et al. (2023). A text cube is an inherently explainable multidimensional structure that allocates documents into cubes along various human-defined dimensions, such as location, date, event, and theme. Each dimension is populated with fine-grained labels (e.g., ‘Miami Downtown’, ‘2021’, ‘Hurricane Irma’, ‘flooding’) extracted from documents. Relevant documents can be easily indexed by these labels represented in cubes. Furthermore, the explicit theme dimension makes the text cube particularly well-suited for theme-specific applications where users Tao et al. (2018) primarily focus on fine-grained inquiries, e.g., where, when, and which hurricane events hit or how much rainfall hurricanes result in certain areas.
@@
In this work, we implement a multi-dimensional (cube) structure called Hypercube, which indexes documents to cubes based on fine-grained labels in a human-defined, multi-dimensional space. Building on this structure, we propose Hypercube-RAG, a novel RAG framework to enhance the retrieval process by retrieving relevant documents from the corresponding cube cells. Hypercube-RAG exhibits three characteristics simultaneously: (1) Efficiency. The retrieval process quickly narrows down the search to the right cube cell(s) based on the labels with the constant time , while taking to fetch documents assigned in that cell. (2) Explainability. Hypercube is inherently explainable as fine-grained labels in cubes represent the compact information in documents (see Figure 2, right and Figure 3); (3) Accuracy. The cube label-based search supports both sparse and dense embedding strategies, effective for capturing uncommon thematic terminology and semantically relevant information (Section 3.3). Overall, our contributions are summarized as follows:
@@
We identify the shortcomings of conventional RAG methods, especially for theme-specific question-answering in scientific domains.
@@
We propose Hypercube-RAG, a simple yet accurate, efficient, and explainable RAG method.
We conduct experiments on three datasets, demonstrating that our Hypercube-RAG outperforms other baseline methods in accuracy and efficiency when it comes to in-domain question-answering.
2 Related Work
2.1 Text-based RAG
@@
BM25 Robertson and Walker (1994); Trotman et al. (2014) is a classical sparse retrieval method based on TF-IDF principles, widely used in information retrieval. It ranks documents by scoring query term matches, with adjustments for term frequency saturation and document length. While efficient and interpretable, BM25 lacks semantic understanding and often misses paraphrased or contextually related documents without exact lexical overlap, limiting its effectiveness in tasks requiring deeper language comprehension. Dense embedding retrievers that compute contextual similarity between the given query and a document/chunk in the embedding space Jiang et al. (2023). Common retrievers consist of DPR Karpukhin et al. (2020), Contriever Izacard et al. (2022), e5 Wang et al. (2022), and ANCE Xiong et al. (2021). By capturing deeper semantic relationships, these methods improve performance in open-domain question answering and generative tasks, especially when surface-form matching fails. However, dense retrieval can return semantically close yet topically irrelevant documents, introducing noise and hallucinations. Moreover, its opaque nature also hinders interpretability Ji et al. (2019), limiting adoption in high-stakes or scientific domains. Recent efforts aim to mitigate these limitations by incorporating hybrid retrieval techniques Maillard et al. (2021) and domain-adaptive retrievers Lee et al. (2021).
2.2 Structured RAG
@@
Graph-based RAG methods enhance the standard RAG paradigm by introducing structured knowledge representations, such as entity or document graphs. GraphRAG Edge et al. (2024) derives an entity knowledge graph from the source documents and gathers summaries for all groups of closely related entities. LightRAG Guo et al. (2024) employs a dual-level retrieval system that enhances comprehensive information retrieval from both low-level and high-level knowledge discovery. HippoRAG Gutiérrez et al. (2024) is a knowledge graph-based retrieval framework inspired by the hippocampal indexing theory of human long-term memory. HippoRAG 2 Gutiérrez et al. (2025) builds on top of HippoRAG and enhances it with deeper passage integration and more effective online use of an LLM. SELF-RAG Asai et al. (2023) enhances an LLM’s quality and factuality through retrieval and self-reflection. RA-DIT Lin et al. (2023) fine-tunes the pre-trained LLMs or retrievers to incorporate more up-to-date and relevant knowledge. Despite these improvements, graph construction and reasoning can be computationally intensive due to massive nodes and edges.
Text Cube.
@@
Wang et al. (2023) designs the text cube to structure spatial-related data across geospatial dimensions, thematic categories, and diverse application semantics. STREAMCUBE implements the text cube to incorporate both spatial and temporal hierarchies Feng et al. (2015). Doc2Cube Tao et al. (2018) automates the allocation of documents into a text cube to support multidimensional text analytics. Despite these elaborate designs, the text cube has not yet been explored in the context of retrieval-augmented generation (RAG) systems.
3 Hypercube-RAG
@@
Figure 4 illustrates the Hypercube-RAG framework through a scientific question-answering example. In the following, we define the hypercube structure (Section 3.1), describe the process for constructing the hypercube (Section 3.2), explain how two sparse and dense embedding retrieval strategies are combined inside of hypercube (Section 3.3), and present how to rank documents retrieved using hypercube (Section 3.4).
3.1 Hypercube Formulation
@@
Given a text corpus with documents: , the hypercube is designed as a multidimensional data structure, , where is the number of hypercube dimensions, and is the dimension. For any document , a text classification algorithm (e.g, TeleClass Zhang et al. (2025)) will assign it to one or more cube cells. This process is equivalent to assigning a -dimensional labels () for a document, , where label represents the category of in dimension .
3.2 Hypercube Construction
For most scientific applications, theme, location, date, event, and/or person are often of high interest. The core idea of the hypercube is to identify fine-grained entities or topics contained in the corpus and index documents in cubes with the appropriate dimensions, thereby facilitating the subsequent retrieval. In this work, we employ multiple methods to construct a holistic and precise knowledge hypercube.
Name Entity Recognition (NER).
We first pre-define hypercube dimensions LOCATION, DATE, EVENT, PERSON and ORGANIZATION. For each dimension, we utilize the pretrained language models (PLMs) Devlin et al. (2019); Liu et al. (2019) to extract salient entities from documents: , where is the set of named entities identified in the document, , and refers to a PLM. For instance, those entities belonging to the EVENT dimension are represented:
Theme-specific Extraction.
While the above entities can be captured by NER models, abstract and intricate theme-specific information may be missed, such as ‘hurricane path’, ‘rainfall intensity’, and ‘climate change’. Therefore, we add another hypercube dimension THEME, such key theme-specific phrases can be extracted. We implement it with KeyBERT111https://github.com/MaartenGr/KeyBERT?tab=readme-ov-file, a language model pre-trained on large-scale text corpora, which demonstrates strong performance in capturing general topical relevance.
Using the methods described above, we extract fine-grained entities, topics, and phrases, and assign them to different hypercube dimensions, including LOCATION, DATE, EVENT, ORGANIZATION, PERSON, and THEME. To enhance coverage of in-domain semantics, we incorporate human-in-the-loop curation to expand the THEME dimension, ensuring it captures a comprehensive set of relevant topics. Additional implementation details are provided in Appendix E.
3.3 Hypercube Retrieval
@@
A constructed knowledge hypercube can serve for the retrieval process based on its fine-grained labels. Given a query , an LLM first decomposes it into a set of entities and topics, , aligned with various dimensions (see the prompt in Figure 4), then retrieval inside of hypercube is performed by matching these decomposed components with cube labels. Our hypercube supports two matching strategies: sparse exact lexical match and dense embedding match. We prioritize leveraging the exact matching strategy due to its high precision, scoring documents based on the frequency with which query entities appear within them:
@@
To address such cases, Hypercube-RAG supports semantic retrieval by computing the similarity between query components and fine-grained cube labels. When the similarity score exceeds a predefined threshold , semantic retrieval is triggered. Specifically, both the query entities and cube labels are projected into an embedding space, , with an encoder. It can be represented as:
@@
3.4 Hypercube Ranking
The retrieved documents need to be provided to the LLM as contextual input. However, long-context RAG methods do not always improve the quality of LLM responses, since they may introduce irrelevant or noisy information Liu et al. (2024); Jin et al. (2025). This raises a critical challenge: how to precisely filter and select theme-relevant documents. To address this, we prioritize returning those documents that fully cover key components derived from the query:
@@ 3.4 Hypercube Ranking
4 Experiments
4.1 Datasets
In this work, we focus on studying Hypercube-RAG within the context of a single hypercube configuration. While more general domains may benefit from modeling multiple hypercubes, our evaluation is centered on one scientific domain to enable a focused and controlled analysis. We also provide a more detailed discussion in Appendix C. We collected the text corpus from a subset of the Science Daily Climate Change dataset (SciDCC) Mishra and Mittal (2021), which was created by web scraping from the “Geography” and “Climate” topics in the environmental science section of the Science Daily website222https://www.sciencedaily.com/. We created an aging dam corpus by collecting news articles related to dam failures in the United States from Google News333https://news.google.com/. Then, an LLM (GPT-4o) is employed to generate 300, 300, and 186 QA pairs, respectively. Each pair was manually validated to ensure that the corresponding question could be answered using information contained in the corpus. To ensure representativeness, each dataset generates three types of QA pairs: (a) long-form, free-style; (b) short-term factual; and (c) quantitatively factual. We present the summary in Table 1 and more details in Appendix B.
| Dataset | # Doc. | Doc. Length | # QA | Ques. Length | Ans. Length |
| Hurricane | 844 | 521781 | 300 | 627 | 2119 |
| Geography | 432 | 221646 | 300 | 534 | 279 |
| Aging Dam | 186 | 254782 | 186 | 633 | 2150 |
4.2 Baselines

### v2

# Hypercube-Based Retrieval-Augmented Generation for Scientific Question-Answering

@@ Abstract.
Large language models (LLMs) often need to incorporate external knowledge to solve theme-specific problems. Retrieval-augmented generation (RAG) has shown its high promise, empowering LLMs to generate more qualified responses with retrieved external data and knowledge. However, most RAG methods retrieve relevant documents based on either sparse or dense retrieval methods or their combinations, which overlooks the essential, multi-dimensional, and structured semantic information present in documents. This structured information plays a critical role in finding concise yet highly relevant information for domain knowledge-intensive tasks, such as scientific question-answering (QA). In this work, we introduce a multi-dimensional (cube) structure, Hypercube, which can index and allocate documents in a pre-defined multi-dimensional space. Built on the hypercube, we further propose Hypercube-RAG, a novel RAG framework for precise and efficient retrieval. Given a query, Hypercube-RAG first decomposes it based on its entities, phrases, and topics along with pre-defined hypercube dimensions, and then retrieves relevant documents from cubes by aligning these decomposed components with corresponding dimensions. Experiments on three datasets across different domains demonstrate that our method improves response accuracy by and retrieval accuracy by over the strongest RAG baseline. It also boosts retrieval efficiency (speed) by one or two magnitudes faster than graph-based RAG. Notably, our Hypercube-RAG inherently offers explainability by revealing those underlying dimensions used for retrieval. The data and code are available at https://github.com/JimengShi/Hypercube-RAG.
1. Introduction
Hypecube- vs. Semantic RAG: A case study
@@ 1. Introduction
Large language models (LLMs) often suffer from hallucinations and factual inaccuracies, especially for scientific question-answering tasks (Augenstein et al., 2024). To address this phenomenon, retrieval-augmented generation (RAG) has emerged as the de facto approach, incorporating external domain knowledge to generate contextually relevant responses (Huang et al., 2025). Despite recent advancements, they exhibit key limitations in theme-specific applications on which the fine-grained topics are highly focused. For instance, retrieved documents are often semantically similar to the query but overlook specialized themes (e.g., terminology or nuanced contextual cues) prevalent in specific literature or reports. Moreover, the retrieval process often suffers from inefficiency and limited transparency, which are critical in high-stakes domains, such as healthcare (Pujari et al., 2023), chemistry (Wellawatte et al., 2025), environment (Wang et al., 2025), and law (Pipitone and Alami, 2024).
@@
The conventional sparse lexical retrieval ranks a set of documents based on the appearance of query terms (Kadhim, 2019). Due to its reliance on the exact token overlap, it precisely captures specific themes or topics, and the retrieval process is efficient and interpretable. However, it struggles to retrieve contextually related or paraphrased documents that require semantic understanding. To mitigate the above limitation, dense embedding retrievers were proposed by computing the similarity between vector embeddings of query-document pairs (Karpukhin et al., 2020). While contextual understanding is relatively improved, the hurdle of missing specific themes remains, causing semantically similar but off-topic retrievals (Krishna et al., 2021; Kang et al., 2024). Figure 1 shows an example where the Semantic-RAG returns documents related to (or the lack of) precipitation in Florida, but overlooks the critical location-specific information for “Melbourne Beach”. Moreover, retrieval in an embedding space makes it challenging to interpret why certain documents were selected (Ji et al., 2019). Additionally, a graph-based RAG represents documents as graph structures, where nodes correspond to entities or concepts and edges capture their relationships (Peng et al., 2024). They offer explanations of the retrieval process by outputting the traversed subgraph. However, two limitations persist still. First, traversing such subgraphs introduces a significant risk of information overload, as the expanding neighborhoods may include substantial amounts of irrelevant themes (Wang et al., 2024a). Second, graph-based retrieval methods are prone to inefficiencies and considerable scalability bottlenecks (Devezas and Nunes, 2021; Zhang et al., 2025a; Han et al., 2024), primarily stemming from the vast number of nodes and edges in their underlying graph structures.
@@
The methods discussed above tend to perform poorly in at least one of the aspects among accuracy, efficiency, and explainability. This motivates us to develop a RAG system that overcomes this weakness. In this pursuit, we identified the text cube as a promising technique (Tao et al., 2018; Wang et al., 2023). A text cube is an inherently explainable multidimensional structure that allocates documents into cubes along various pre-defined dimensions, such as location, date, event, and specific-theme(s). Each cell is associated with fine-grained entities (e.g., “Miami Downtown”, “2021”, “Hurricane Irma”, “flooding”) extracted from documents, which are viewed as document labels. Relevant documents can be indexed and located in cube cells associated with these document labels (cube labels for brevity). Moreover, the theme dimension specifies the themes in hierarchies which make the text cube well-suited for theme-specific applications where users primarily focus on fine-grained inquiries (Tao et al., 2018), such as which hurricane hits and how much rainfall it results in certain areas, or which medicine is effective in treating which type of cancer.
@@
In this work, we implement a multi-dimensional (cube) structure called Hypercube, which indexes documents to cube cells based on their fine-grained labels in a pre-defined, multi-dimensional semantic space. Building on this cube structure, we further propose Hypercube-RAG, a novel RAG framework to enhance the retrieval process by retrieving relevant documents from the corresponding cube cells. Compared with the aforementioned RAG approaches, our Hypercube-RAG simultaneously exhibits three critical characteristics due to its multidimensionality: (1) High Accuracy. The multidimensional, cube label-based search supports both sparse and dense embedding strategies, effective for capturing essential information in multi-dimensional space as well as uncommon thematic terminology and semantically relevant cases. (2) High Efficiency. The retrieval mechanism quickly narrows down the search to the pertinent cube cell(s) based on their associated labels. This operation exhibits a constant time complexity of , which could be multiple magnitudes faster, thereby demonstrating potential robust scalability concerning corpus size. (3) Explainability. Hypercube is inherently explainable as fine-grained labels associated in cube cells represent the compact information in documents (see Figure 2 (right) and Figure 4). Overall, our contributions are as follows:
@@
We identify the shortcomings of conventional RAG approaches, especially for theme-specific question-answering tasks in the scientific domains.
@@
We design a multi-dimensional structure (hypercube) and further propose Hypercube-RAG, a simple yet accurate, efficient, and explainable RAG method.
We conduct experiments on three datasets across different domains with multiple metrics, demonstrating that our approach outperforms other baseline methods consistently.
2. Related Work
2.1. Text-based RAG
@@
BM25 (Robertson and Walker, 1994; Trotman et al., 2014) is a classical sparse retrieval method based on TF-IDF principles, widely used in information retrieval. It ranks documents by scoring query term matches, with adjustments for term frequency saturation and document length. While efficient and interpretable, BM25 lacks semantic understanding and often misses paraphrased or contextually related documents without exact lexical overlap, limiting its effectiveness in tasks requiring deeper language comprehension. Dense embedding retrievers compute contextual similarity between the given query and a document/chunk in the embedding space. Common retrievers consist of DPR (Karpukhin et al., 2020), Contriever (Izacard et al., 2022), e5 (Wang et al., 2022), and ANCE (Xiong et al., 2021). By capturing deeper semantic relationships, these methods improve performance in open-domain question answering and generative tasks, especially when surface-form matching fails. However, dense retrieval can return semantically close yet topically irrelevant documents, introducing noise and hallucinations. Moreover, its opaque nature also hinders interpretability (Ji et al., 2019), limiting adoption in high-stakes or scientific domains. Recent efforts aim to mitigate these limitations by incorporating hybrid retrieval techniques (Maillard et al., 2021) and domain-adaptive retrievers (Lee et al., 2021).
2.2. Structured RAG
@@
Graph-based RAG methods enhance the standard RAG paradigm by introducing structured knowledge representations, such as entity or document graphs. GraphRAG (Edge et al., 2024) derives an entity knowledge graph from the source documents and gathers summaries for all groups of closely related entities. LightRAG (Guo et al., 2024) employs a dual-level retrieval system that enhances comprehensive information retrieval from both low-level and high-level knowledge discovery. HippoRAG (Gutiérrez et al., 2024) is a knowledge graph-based retrieval framework inspired by the hippocampal indexing theory of human long-term memory. HippoRAG 2 (Gutiérrez et al., 2025) builds on top of HippoRAG and enhances it with deeper passage integration and more effective online use of an LLM. SELF-RAG (Asai et al., 2023) enhances an LLM’s quality and factuality through retrieval and self-reflection. RA-DIT (Lin et al., 2023) fine-tunes the pre-trained LLMs or retrievers to incorporate more up-to-date and relevant knowledge. Despite these improvements, graph construction and reasoning can be computationally intensive due to the vast number of nodes and edges.
@@
Text Cube. The text cube has been proposed to structure spatial-related data across geospatial dimensions, thematic categories, and diverse application semantics (Wang et al., 2023). STREAMCUBE implements the text cube to incorporate both spatial and temporal hierarchies (Feng et al., 2015). Doc2Cube (Tao et al., 2018) automates the allocation of documents into a text cube to support multidimensional text analytics. Despite these elaborate designs, the text cube has not yet been explored in the context of retrieval-augmented generation (RAG) systems.
3. Our Method: Hypercube-RAG
@@
Figure 3 illustrates the Hypercube-RAG framework through a concrete question-answering example. In the following, we define the hypercube structure (Sec. 3.1), describe the process for designing and constructing the hypercube (Sec. 3.2), explain how two sparse and dense embedding retrieval strategies are combined inside of hypercube (Sec. 3.3), and present how to rank retrieved documents (Sec. 3.4).
3.1. Hypercube Formulation
@@
Given a text corpus with documents , the hypercube is designed as a multidimensional structure, , where is the number of hypercube dimensions, and is the dimension. For any document , a text classification algorithm (e.g, TeleClass (Zhang et al., 2025b)) will assign it to one or more cube cells. This process is equivalent to assigning a -dimensional labels () for a document, where label represents a category of document in dimension .
3.2. Hypercube Design and Construction
Hypercube Design. The core idea of the hypercube is to identify fine-grained entities or topics present in the corpus and to index documents into cube cells along appropriate dimensions, thereby enabling the subsequent retrieval. In this work, we design the hypercube structure by systematically determining its dimensions. First, we employ pre-trained language models to extract entities from the entire corpus. Next, the K-means clustering algorithm is applied to group these entities into semantically coherent clusters. For each cluster, we let a large language model (LLM) generate concise summaries that characterize the underlying cluster categories or dimensions111For example, LLM outputted “these entities in the cluster include countries, states, provinces, cities, towns, and specific addresses.”. Finally, we consolidate the candidate dimensions across clusters to construct a high-level, multi-dimensional space for document indexing.
This dimension-identification process for hypercube design has demonstrated strong alignment with domain expert expectations and yielded promising performance (although diverse strategies remain open for future exploration).
In Appendix A, we provide a detailed dimension-identification process for each dataset.
Hypercube Construction. With the hypercube dimensions finalized or pre-defined, we refine the previously generated concise summaries as prompts to an LLM to extract entities associated with each high-level dimension. These extracted entities are then used as labels to index documents into their corresponding cube cells. The process to extract entities for one dimension is represented as:
where represents the extracted entity set from a document, , along the dimension, , refers to an LLM, and denotes the corresponding prompt used for . We present a prompt example in Appendix B.
3.3. Hypercube Retrieval
@@
A constructed knowledge hypercube can serve for the retrieval process based on its fine-grained labels. Given a query , an LLM first decomposes it into a set of entities, , aligned with pre-defined dimensions (see the prompt in Figure 3), then retrieval inside of hypercube is performed by matching these decomposed components with labels associated in cube cell(s). Our hypercube supports two matching strategies: sparse exact lexical match and dense embedding match. We prioritize leveraging the exact matching strategy due to its high precision, scoring documents based on the frequency with which query entities appear:
@@
To address such cases, Hypercube-RAG supports semantic retrieval by computing the similarity between query components and fine-grained cube labels, along with a certain dimension, . When the similarity score exceeds a predefined threshold , semantic retrieval is triggered. Specifically, both the query entities and cube labels are projected into an embedding space, , with an encoder. It can be represented as:
@@
3.4. Hypercube Ranking
The retrieved documents need to be provided to the LLM as contextual input. However, long-context RAG methods do not always improve the quality of LLM responses, since they may introduce irrelevant or noisy information (Liu et al., 2024; Jin et al., 2025). This raises a critical challenge: how to precisely filter and select theme-relevant documents. To address this, we prioritize returning those documents that fully cover key components derived from the query:
@@ 3.4. Hypercube Ranking
4. Experiments
4.1. Datasets
To verify the applicability of Hypercube-RAG, three datasets over varying types across different domains are selected: (1) SciFact (Wadden et al., 2020): a dataset of expert-written scientific claims paired with evidence-containing biomedical abstracts annotated with labels and rationales. (2) LegalBench (Pipitone and Alami, 2024): a dedicated QA and retrieval benchmark in the legal domain, entirely human-annotated by legal experts. (3) SciDCC (Mishra and Mittal, 2021): a synthesized dataset using LLMs in the geo-science domain, covering “Geography”, “Hurricane”, and “aging dams”. We collected the text corpus from a subset of the Science Daily Climate Change dataset (SciDCC) (Mishra and Mittal, 2021) from the Science Daily website222https://www.sciencedaily.com/. “Aging dams” discusses dam failures in the United States, whose corpus comprises related news articles downloaded from Google News333https://news.google.com/, collected by domain experts. The summary is presented in Table 1.
| Datasets | # Docs | # QA | QA type | Ques. Length | Ans. Length |
| SciFact (Medical) | 5813 | 188 | True or False | 528 | 1 |
| LegalBench (Law) | 95 | 977 | Open-Ended | 1639 | 10304 |
| SciDCC (Geoscience) | 1462 | 789 | Open-Ended | 636 | 10150 |
In this work, we focus on studying Hypercube-RAG with one hypercube for each dataset, while more complex domains across various themes may benefit from multiple hypercubes. We provide a detailed discussion on multiple hypercubes in Appendix C.
| Method | SciFact (Medical) | LegalBench (Law) | SciDCC (Geoscience) | |||||||||
[…]
| F1 | Semantic | Correctness | Completeness | F1 | Semantic | Correctness | Completeness | F1 | Semantic | Correctness | Completeness | |
[…]
| GPT-4o | 81.3 | 84.9 | 81.9 | 81.9 | 22.6 | 56.8 | 52.7 | 32.1 | 26.5 | 64.6 | 59.8 | 29.7 |
[…]
| BM25 (Robertson and Walker, 1994) | 87.8 | 89.8 | 87.7 | 87.7 | 31.8 | 61.2 | 54.6 | 50.9 | 45.3 | 76.6 | 79.1 | 68.2 |
| Contriever (Izacard et al., 2021) | 72.3 | 77.1 | 72.3 | 72.3 | 37.8 | 62.4 | 67.1 | 64.3 | 42.6 | 76.5 | 77.2 | 65.6 |
| e5 (Wang et al., 2024b) | 90.4 | 92.1 | 90.4 | 90.4 | 38.3 | 63.2 | 74.1 | 71.5 | 45.0 | 77.3 | 81.9 | 69.4 |
| NV-Embed-v2 (Lee et al., 2025) | 91.2 | 92.6 | 91.3 | 91.3 | 32.8 | 60.7 | 54.1 | 50.3 | 44.9 | 77.8 | 80.4 | 69.8 |
[…]
| LightRAG (Guo et al., 2024) | 67.3 | 69.3 | 76.6 | 76.6 | 28.5 | 61.2 | 68.9 | 66.5 | 30.9 | 68.3 | 66.7 | 51.9 |
| GraphRAG (Edge et al., 2024) | 84.9 | 86.2 | 84.8 | 84.8 | 36.3 | 61.8 | 68.4 | 64.7 | 39.2 | 71.4 | 70.4 | 56.6 |
| HippoRAG (Gutiérrez et al., 2024) | 85.2 | 87.3 | 86.1 | 86.1 | 34.7 | 60.4 | 67.5 | 63.5 | 41.6 | 72.4 | 75.4 | 59.1 |
| HippoRAG 2 (Gutiérrez et al., 2025) | 88.3 | 90.2 | 89.3 | 89.3 | 37.2 | 62.6 | 70.9 | 69.3 | 43.1 | 75.1 | 84.7 | 58.9 |
[…]
| Hypercube-RAG | 91.5 | 93.0 | 91.5 | 91.5 | 40.5 | 64.6 | 81.5 | 78.1 | 46.4 | 80.3 | 83.3 | 72.2 |
| Method | SciFact (Medical) | LegalBench (Law) | SciDCC (Geoscience) | |||||||||
[…]
| Precision@3 | Recall@3 | Precision@5 | Recall@5 | Precision@3 | Recall@3 | Precision@5 | Recall@5 | Precision@3 | Recall@3 | Precision@5 | Recall@5 | |
[…]
| BM25 (Robertson and Walker, 1994) | 27.5 | 75.9 | 18.1 | 81.3 | 12.0 | 36.1 | 7.8 | 38.8 | 28.6 | 86.0 | 17.9 | 89.7 |
| Contriever (Izacard et al., 2021) | 13.8 | 35.9 | 8.9 | 38.1 | 19.2 | 57.5 | 12.7 | 63.3 | 17.8 | 53.2 | 12.7 | 62.4 |
| e5 (Wang et al., 2024b) | 33.0 | 87.7 | 21.2 | 92.8 | 21.7 | 65.1 | 13.9 | 69.5 | 29.0 | 87.1 | 18.4 | 91.2 |
| NV-Embed-v2 (Lee et al., 2025) | 34.4 | 87.9 | 22.9 | 95.2 | 13.4 | 40.3 | 8.7 | 43.6 | 19.7 | 59.4 | 12.5 | 62.4 |
| HippoRAG 2 (Gutiérrez et al., 2025) | 34.7 | 88.2 | 22.6 | 95.6 | 12.8 | 38.8 | 12.7 | 64.3 | 18.2 | 57.6 | 12.1 | 61.7 |
[…]
| Hypercube-RAG | 37.5 | 88.8 | 27.6 | 96.1 | 26.1 | 75.4 | 16.9 | 79.7 | 35.5 | 86.0 | 28.8 | 92.4 |
4.2. Baselines